# H-DES: a Quantum-Classical Hybrid Differential Equation Solver

### v1

@@ Abstract
In this article, we introduce an original hybrid quantum-classical algorithm based on a variational quantum algorithm for solving systems of differential equations. The algorithm relies on a spectral method, which involves encoding the solution functions in the amplitudes of the quantum states generated by different parametrized circuits and transforms the task of solving the differential equations into an optimization problem. We first describe the principle of the algorithm from a theoretical point of view. We provide a detailed pseudo-code of the algorithm, on which we conduct a complexity analysis to highlight its scaling properties. We apply it to a set of examples, showcasing its applicability across diverse sets of differential equations. We discuss the advantages of our method and potential avenues for further exploration and refinement. [^1] The algorithm described in this paper is subject to a patent submission (EP24306601). Use of this method requires a license, which can be obtained from ColibriTD.
@@ 1 Introduction
In this paper, we propose a variational quantum algorithm to solve partial DEs based on a spectral decomposition of the solution function. This approach avoids the problem of requiring an exponentially growing number of circuits to evaluate the derivative of the solution function but requires the efficient readout of expectation values. This article is organized as follows. In Sec. 2, we introduce the algorithm and its key principle and detail the steps of its workflow. In Sec. 3, we discuss the application of the algorithm to multivariate equations. In Sec. 4, we explain some theoretical elements concerning the complexity and scaling of the hybrid algorithm. In Sec. 5, we run the algorithm on several systems of DEs, showcasing the versatility of our method. In Sec. 6, we discuss the advantages of our algorithm and compare it to other approaches in the literature. In Sec. 7, we conclude the article and present possible improvements of our algorithm. In the appendices, we share how one can deal with boundary conditions in simple cases (see App. A), some examples to illustrate the theoretical ideas behind the algorithm (see App. B), more details from the complexity study (see App. C), and a theoretical framework for performance comparisons of several solvers (see App. D).
@@ 2.1 Variational Quantum Circuit and Ansatz
To construct the parametrized state used to minimize the loss function, we need a variational quantum circuit (VQC). The architectures employed here are based on the structure of the hardware-efficient Ansatz (HEA). In the HEA, each set of parametrized single-qubit rotations is followed by an entangling layer of CNOT gates. This block of rotations and CNOTs is then repeated $d$ times, where $d$ is called the depth of the VQC. In what follows, we consider linear connectivity among qubits, but the CNOTs can be easily generalized to other connectivity architectures [LOCC24].
@@
Generally, the rotation layer includes rotations around different axes, i.e., $R_{x}$, $R_{y}$, and $R_{z}$ gates, but in our case, this layer is composed only of parametrized $R_{y}$ gates for each qubit (see Fig. 2). This choice was driven by the fact that we do not need to generate states with complex amplitudes since only the associated probabilities are used to encode the solution function. We thus chose to restrict the research space to real states, for which using only $R_{y}$ gates is sufficient. We also empirically observed that results for different loss functions reached satisfactory precision with a single rotation gate per layer (at least in the ideal case, i.e., with perfect operations in the circuit).
[…]
@@
2.2 Spectral methods and Chebyshev polynomials
Our approach is based on the spectral method, in which the solution function is expressed as a linear combination of the elements of a complete basis of orthogonal functions. Therefore, the task of finding the best solution to a DE consists of determining the best approximation of the coefficients of the solution in that basis.
We choose to express the solution function in the basis of Chebyshev polynomials, which are widely used as basis functions for spectral methods [Boy01, WS07]. These polynomials form a complete orthogonal system and allow us to represent piecewise smooth and continuous functions in finite domains. The expression of their derivatives is straightforward and can also be defined by recurrence, making it easy and efficient to implement.
@@ 2.3 Encoding and evaluation of the function
In this subsection, we detail how we use the $n$-qubit state generated by the variational quantum circuit to represent the solution function. We denote as |$\ket{\psi_{f}}$ the state coming out of the VQC modeled by the gate $U_{\theta}$, expressed in the computational basis (with decimal notation) as follows:
|$\ket{\psi_{f}}=U_{\theta}\ket{0}^{\otimes n}=\sum_{i=0}^{2^{n}-1}a_{i}\ket{i}~{},$ (3)
with $\text{with }~{}~{}\sum_{i=0}^{2^{n}-1}|a_{i}|^{2}=1~{}.$ (4)
@@
Since probabilities are positive real numbers, expressing $f$ using only positive coefficients in the basis would not allow us to represent all continuous functions. One would also need to combine positive and negative coefficients in the decomposition of the solution function. In order to do this, we add to the list of polynomial values a copy of each one, multiplied by $-1$. In the quantum encoding of our function, these probabilities are associated with the second half of the basis states (see Eq. 7).
Furthermore, in the particular case $x\in[-1,1]$, we have $\text{Cheb}(k,x)\in[-1,1]$ for all $x$ and $k$. When we take the linear combination of Chebyshev polynomials with the associated difference of probabilities (because of the duplication of Chebyshev polynomials with a minus sign), which are bounded by the normalization constrain, the resulting function is also bounded. Therefore, to allow for the modeling of any general solution function, we introduce a scaling parameter $\lambda\in\mathbb{R}$.
@@
Therefore, the variational circuit, for a given set of parameters $\theta=\{\theta_{i}\}_{i}$, models the state that encodes the decomposition of the function in the Chebyshev polynomial basis. Note that in the case of a system of DEs, each solution function has a different circuit and one can choose the number of qubits to allocate for each circuit independently.
@@
$\begin{gathered}O_{C}(x)=\sigma_{z}\otimes\left(\sum_{i=0}^{2^{n-1}-1}\text{Cheb}(i,x)\outerproduct{i}{i}\right)\\ =\begin{pmatrix}\text{Cheb}(0,x)\\ &\ddots&&&&\text{\huge 0}\\ &&\text{Cheb}(2^{n-1}-1,x)\\ &&&-\text{Cheb}(0,x)\\ \text{\huge 0}&&&&\ddots\\ &&&&&-\text{Cheb}(2^{n-1}-1,x)\\ \end{pmatrix}~{}.\end{gathered}$ (7)
@@ 2.4 Evaluation of the derivatives
One of the most remarkable advantages of this algorithm is the simplicity with which any derivative of a function can be calculated, and this is due to two characteristics of this method. The first is that a derivative of any order can expressed as
@@
$𝜓𝑓\frac{\partial^{q}f(x)}{\partial x^{q}}=\lambda~{}\expectationvalue{O_{\partial^{q}C}(x)}{\psi_{f}}~{},$ (9)
@@ 2.5 Loss function
In order to build the loss-function, we collect all the terms of the DE on the left side, i.e., $E_{i}(x)=0$, where $E$ is the set of differential equation in our system, and we thus consider the $i^{th}$ differential equation. The evaluation of the DEs requires the evaluation of all the functions and their derivatives (appearing at least once) at each sample point, which in turn implies the computation of expectation values of the right observables.
@@
In the case of a linear DE involving only one function, one can evaluate it for a specific point $x_{s}$ by evaluating only one expectation value. In fact, the DE is defined as the linear combination of the solution function, its derivatives, and potentially the sample point itself. These terms are evaluated by computing the expectation value of an observable ($O_{\partial^{q}C}(x)$ for its $q^{th}$ derivative, and the scaled identity $x_{s}I$ for the sample point) on the same state. Therefore, the DE can be evaluated by regrouping all the observables into a single one using the linearity of the inner product.
@@
where the first term evaluates the match with respect to the DEs, while the second accounts for the boundary conditions.
@@
To construct both terms, we use the mean squared error (MSE). Using MSE metrics, the differential part of the loss function can be explicitly written as
@@
where the $x_{s}$ are the coordinates of the sample points, $\theta$ is the set of angles parametrizing the circuit. Similarly, the boundaries component is computed as the sum of all errors for each boundary condition for all functions appearing in the DE. This can be expressed as follows:
@@ 2.6 Validation of the results
We thus obtain a validation score that can be used to compare solutions for different sets of DEs, different solver parameters, or different solvers entirely. The genericity of the validation score allows us to then establish an acceptance threshold below which we consider the set of DEs solved. This can be dictated by the precision needed in different fields or industries, or for instance by benchmarking several sets of DEs and tuning the threshold to match a given acceptance criteria. In App. D, we present a theoretical framework to describe and compare the quality of convergence of several solvers, within which the global validation score $V$ can be used to measure performance.
@@ 3 Multivariate functions
In this subsection, we detail the multivariate case, i.e., when a solution function $f$ takes as input $X=(x_{1},x_{2},\dots,x_{v})\in\mathbb{R}^{v}$. Since the variable dependence is encoded only in the observables, moving from 1-dimensional to $v$-dimensional space does not affect the rest of the algorithm.
In the uni-dimensional case, each basis state (of the computational basis) is used to represent, in binary notation, the order of a Chebyshev polynomial. Except for the first qubit, which is used to determine the sign in front of the polynomial, all the remaining qubits of the basis states are indeed used to encode the corresponding order. In the multivariate case, we assign to each variable $x_{j}$ a specific number of qubits that are used to encode the order of the corresponding Chebyshev polynomial. In fact, one possible way of defining the multivariate Chebyshev polynomial $\text{Cheb}(i,x_{1},x_{2},\dots,x_{v})$ is to take the product of independent univariate Chebyshev polynomials $\text{Cheb}(L_{j},x_{j})$ defined on a rectangular domain [PV11, Tre16, Mas80].
Therefore, each basis state |$\ket{i}$ of the $(n+1)$-qubit state modeling the solution function is used to encode the different indices assigned to each variable. We write the binary decomposition of the basis state |$\ket{i}$ (written in decimal notation) as
@@
For each variable $x_{j}$, we attribute $l_{j}$ qubits to encode the order (of the Chebyshev polynomial) and denote as $L_{j}$ the binary decomposition of the order. This allows us to separate the basis state |$\ket{i}$ as the tensor product of the first (sign) qubit with $v$ states, such that
@@
$O_{C}(x_{1},x_{2},\dots,x_{v})=\sigma_{z}\otimes\left(\sum_{i=0}^{2^{n-1}-1}\text{Cheb}(i,x_{1},x_{2},\dots,x_{v})\outerproduct{i}{i}\right)~{}.$
@@ 3.1 Computing the derivatives efficiently
When we move to the case of second-order derivatives ($g=2$), we need to distinguish between two cases: that of a partial derivative that is second order in a single variable and that of a partial derivative in two distinct variables. In the first case, we only need to elevate the derivative to second order as follows:
@@
[…]
@@
partial derivatives, which correspond to the sum of all possible partial derivatives at each order, with $n_{d}=\max(G_{F})$ the highest derivative order involved in the DEs and $n_{v}$ the maximum number of variables in functions involved in the DEs (see next section for the definition of $G_{F}$). Consequently, the computational cost of computing all the partial derivatives, expressed as a function of the cost $C_{b}$ of a multiplication/division, is equal to:
@@
To compute all these partial derivatives, we only need to evaluate and multiply the value of the $d^{th}$ partial derivatives in one variable $\frac{\partial^{d}\text{Cheb}}{\partial{x_{j}}^{d}}(L_{j},x_{j})$. For each value of $d$, we have $n_{v}$ such partial derivatives to compute. In the worst case, we have to prepare all the possible partial derivatives of that form. This means that, for half of the basis states (since the other half only introduce a minus sign in front of the Chebyshev polynomials) and all sample points, in the worst case, we compute and store $2^{n-1}\times n_{v}\times n_{d}\times n_{s}$ values.

### v2

# affiliationtext: ColibriTD, 91 Rue du Faubourg Saint Honoré, 75008 Paris, France

@@ Abstract
In this article, we introduce an original hybrid quantum-classical algorithm based on a variational quantum algorithm for solving systems of differential equations. The algorithm relies on a spectral decomposition of the trial functions that are encoded directly in the quantum states generated by different parametrized circuits, and transforms the task of solving the differential equations into an optimization problem. We first describe the principle of the algorithm from a theoretical point of view. We provide a detailed pseudo-code of the algorithm, on which we elaborate preliminary elements for a complexity analysis to highlight some of its scaling properties. We apply our algorithm to a set of examples, running on emulators and real hardware showcasing its applicability across diverse sets of differential equations. We discuss the advantages of our method and potential avenues for further exploration and refinement. [^1] The algorithm described in this paper is subject to a patent submission (EP24306601). Use of this method requires a license, which can be obtained from ColibriTD.
@@ 1 Introduction
In this paper, we propose an original variational quantum algorithm to solve partial differential equations (PDEs) based on a spectral decomposition of the solution function. The H-DES algorithm has been designed to run on NISQ devices in order to address industrial simulation demands with the available quantum machines. The goal of the article is to present the general workflow of the algorithm, its flexibility and reliability, and how the related toolbox can be used in practice to solve DEs.
This article is organized as follows. In Sec. 2, we introduce the algorithm and its key principle and detail the steps of its workflow. In Sec. 3, we discuss how the algorithm adapts to the case of multivariate equations. In Sec. 4, we provide some premilinary theoretical elements concerning the complexity and scaling of the hybrid algorithm. In Sec. 5, we run the algorithm on several systems of DEs, showcasing the versatility of our method. In Sec. 6, we discuss the advantages of our algorithm and compare it to other approaches in the literature. In Sec. 7, we conclude the article and present possible improvements of our algorithm. In the appendices, we share how one can deal with boundary conditions in simple cases (see App. A), some examples to illustrate the theoretical ideas behind the algorithm (see App. B), more details from the complexity study (see App. C), and a theoretical framework for performance comparisons of several solvers (see App. LABEL:app:performance_profile).
@@ 2.1 Variational Quantum Circuit and Ansatz
To construct the parametrized state used to minimize the loss function, we need a variational quantum circuit (VQC). The architectures employed here are usually based on the structure of the hardware-efficient Ansatz (HEA), but not limited to it. In typical HEA setup, each set of parametrized single-qubit rotations is followed by an entangling layer of controlled (usually CNOT) gates. This block of rotations and controlled gates is then repeated $d$ times, where $d$ is called the depth of the VQC. In what follows, we consider linear connectivity among qubits, but the controlled gates can be easily generalized to other connectivity architectures [LOCC24].
@@
Generally, the rotation layer includes rotations around different axes, i.e., $R_{x}$, $R_{y}$, and $R_{z}$ gates, but in the case of a basis-state encoding (see Sec. 2.3), a possible choice for this layer is to only use parametrized $R_{y}$ gates for each qubit (see Fig. 2). This choice is driven by the fact that, in the case of basis-state encoding, we do not need to generate states with complex amplitudes since only the associated probabilities are used to encode the solution function. We thus chose to restrict the research space to real states, for which using only $R_{y}$ gates is sufficient. We also empirically observed that results for different loss functions reached satisfactory precision with a single rotation gate per layer (at least in the ideal case, i.e., with perfect operations in the circuit).
@@
When running on real QPUs (Quantum Processing Unit), the circuit has to go through a transpilation process to express the circuit by means of unitary gates supported by the QPU (also called native gates). In that case, building the ansatz directly by using native gates can help to avoid extra-transpilation steps that usually increase the depth of the circuit and thus the risk of gate-noise propagation and decoherence during execution. Taking into account the qubit connectivity directly in the selection of qubits or the design of the ansatz also helps to reduce the number of additional gates (and thus noise) introduced at the transpilation step.
2.2 Spectral decomposition and Chebyshev polynomials
Our approach is based on the spectral decomposition, in which the solution function is expressed as a linear combination of the elements of a complete basis of orthogonal functions. Therefore, the task of finding the best solution to a DE consists of determining the best approximation of the coefficients of the solution in that basis.
One possible choice is to express the solution function in the basis of Chebyshev polynomials, which are widely used as basis functions for spectral methods [Boy01, WS07]. These polynomials form a complete orthogonal system and allow us to represent piecewise smooth and continuous functions in finite domains. The expression of their derivatives is straightforward and can also be defined by recurrence, making it easy and efficient to implement.
@@ 2.2 Spectral decomposition and Chebyshev polynomials
In the H-DES algorithm, the choice of the basis is not only limited to Chebyshev polynomials, but allows one to select any other suitable complete basis (Legendre, Laguerre, Fourier, B-Splines and NURBS, Walsh functions, …), or even modify and combine several bases. This is a possible way to impose exactly some boundary conditions. In the following sections, we will, by default, discuss the principle of the algorithm by considering Chebyshev polynomials as a basis choice.
@@ 2.3 Encoding and evaluation of the function
[…]
$\text{with }~~\sum_{i=0}^{2^{n}-1}|a_{i}|^{2}=1~.$ (4)
2.3.1 Basis-state encoding
@@
Since probabilities are positive real numbers, expressing $f$ using only positive coefficients in the basis would not allow us to represent all continuous functions. One would also need to combine positive and negative coefficients in the decomposition of the solution function. In order to do this, we add to the list of polynomial values a copy of each one, multiplied by 1-1. In the quantum encoding of our function, these probabilities are associated with the second half of the basis states (see Eq. 7).
Furthermore, in the particular case $x\in[-1,1]$, we have $\text{Cheb}(k,x)\in[-1,1]$ for all $x$ and $k$. When we take the linear combination of Chebyshev polynomials with the associated difference of probabilities (because of the duplication of Chebyshev polynomials with a minus sign), which are bounded by the normalization constraint, the resulting function is also bounded. Therefore, to allow for the modeling of any general solution function, we introduce a scaling parameter $\lambda\in\mathbb{R}$.
@@
Therefore, the variational circuit, for a given set of parameters $\theta=\{\theta_{i}\}_{i}$, models the state that encodes the decomposition of the function in the Chebyshev polynomial basis. Note that in the case of a system of DEs, each solution function has a different circuit, and one can choose the number of qubits to allocate for each circuit independently.
@@
[…]
@@
In the following sections, we will, by default, discuss the principle of the algorithm by considering the global and diagonal observables corresponding to basis-state encoding. The number of probabilities to be estimated grows exponentially with the number of qubits. Moreover, the global nature of the observable implies that the cost function is also global, which may lead to the appearance of barren plateaus, which is another significant limitation on the range of applicability. In the next subsection, we introduce an alternative encoding to overcome these issues.
2.3.2 Pauli-monomial encoding
Another way of encoding the spectral basis in the observable is to consider a linear combination of Pauli monomials instead of probabilities of basis states. This provides another general framework to compute the evaluation of the trial function, with more freedom on the number of orthogonal functions and the choice of the Pauli monomial appearing in the decomposition, which will determine the size of the spectral basis. This is particularly interesting when dealing with barren plateaus, measurement strategies, and readout error [cerezo2021, kashif2025].
Therefore, in the case of a Pauli monomial encoding, the observable $O_{C}(x)$ can be expressed as the linear combination of $M\in\mathbb{N}^{*}$ Pauli monomials such that:
$O_{C}(x)=\sum_{i=0}^{M-1}o_{i}\cdot\text{Cheb}(i,x)\cdot P_{i}~,$ (8)
with, for all $i\in\llbracket 0,M-1\rrbracket$, $o_{i}\in\mathbb{R}^{*}$ a ponderation factor and $P_{i}\in\{I,X,Y,Z\}^{\otimes n}$ an element of the set of generalized Pauli operators, also called $n$-qubit Pauli monomial.
A specific subset of the generalized Pauli ensemble of interest can be chosen to define the loss function more locally. Indeed, the $k$-local observables [kitaev2002, peruzzo2014, cubitt2016, cerezo2021] are defined as observables for which the Pauli monomials appearing in their decomposition are acting non-trivially on at most $k$ qubits. A given Pauli monomial $P_{i}=P_{i,1}\otimes P_{i,2}\otimes\cdots\otimes P_{i,n}$, is said to act non-trivally on $k$ qubits (also called $k$-body interaction) if there are exactly nkn-k Pauli atoms $P_{i,j}$ that are the (one-qubit) identity operator.
The maximum possible number $M_{k}$ of Pauli monomials appearing in the decomposition of a general $k$-local observable is given by:
$M_{k}=\sum_{i=1}^{k}3^{i}\binom{n}{i}~$. (9)
@@ 2.4 Evaluation of the derivatives
In the case where we chose Chebyshev polynomials as a basis, one of the most remarkable advantages of this algorithm is the simplicity with which any derivative of a function can be calculated, and this is due to two characteristics of this method. The first characteristic is that a derivative of any order can expressed as
@@
[…]
@@ 2.5 Loss function
In order to build the loss function, we collect all the terms of the DE on the left side, i.e., $E_{i}(x)=0$, where $E$ is the set of differential equation in our system, and we thus consider the $i^{th}$ differential equation. The evaluation of the DEs requires the evaluation of all the functions and their derivatives (appearing at least once) at each sample point, (sometimes called collocation points), which in turn implies the computation of expectation values of the right observables.
@@
In the case of a linear DE involving only one function, one can evaluate it for a specific point $x_{s}$ by evaluating only one expectation value. In fact, the DE is defined as the linear combination of the solution function, its derivatives, and potentially the sample point itself. These terms are evaluated by computing the expectation value of an observable ($O_{\partial^{q}C}(x)$ for its $q^{th}$ derivative, and the scaled identity $x_{s}I$ for the sample point) on the same state. Therefore, the DE can be evaluated by regrouping all the observables into a single one using the linearity of the inner product, which will make the algorithm essentially similar to a Variational Quantum Eigensolver. This is also possible for non-linear DEs, when the Pauli monomials in the decomposition of the observables (for function and derivatives evaluation) are all commuting. In that case, using a Pauli grouping process, it is sufficient to evaluate the expectation value of a single monomial to retrieve the others by post-processing.
@@
where the first term evaluates the residual of the DEs, while the second accounts for the boundary conditions.
@@
To construct both terms, one can use the mean squared error (MSE). Using MSE metrics, the differential part of the loss function can be explicitly written as
@@
where the $x_{s}$ are the coordinates of the sample points, $\theta$ is the set of angles parametrizing the circuit. Similarly, the boundaries component can be computed as the sum of all errors for each boundary condition for all functions appearing in the DE. This can be expressed as follows:
@@ 2.6 Validation of the results
We thus obtain a validation score that can be used to compare solutions for different sets of DEs, different solver parameters, or different solvers entirely. The genericity of the validation score allows us to then establish an acceptance threshold below which we consider the set of DEs solved. This can be dictated by the precision needed in different fields or industries, or, for instance, by benchmarking several sets of DEs and tuning the threshold to match a given acceptance criterion. In App. LABEL:app:performance_profile, we present a theoretical framework to describe and compare the quality of convergence of several solvers, within which the global validation score $V$ can be used to measure performance.
@@ 3 Multivariate functions
In this subsection, we detail the multivariate case, i.e., when a solution function $f$ takes as input $X=(x_{1},x_{2},\dots,x_{v})\in\mathbb{R}^{v}$. Since the variable dependence is encoded only in the observables, moving from 1-dimensional to $v$-dimensional space does not affect the rest of the algorithm. We recall that we focus on the basis-state encoding of the trial function, in the basis of Chebyshev polynomials.
In the unidimensional case, each basis state (of the computational basis) is used to represent, in binary notation, the order of a Chebyshev polynomial. Except for the first qubit, which is used to determine the sign in front of the polynomial, all the remaining qubits of the basis states are indeed used to encode the corresponding order. In the multivariate case, we assign to each variable $x_{j}$ a specific number of qubits that are used to encode the order of the corresponding Chebyshev polynomial. In fact, one possible way of defining the multivariate Chebyshev polynomial $\text{Cheb}(i,x_{1},x_{2},\dots,x_{v})$ is to take the product of independent univariate Chebyshev polynomials $\text{Cheb}(L_{j},x_{j})$ defined on a rectangular domain [PV11, Tre16, Mas80].
@@
[…]
@@
[…]
@@ 3.1 Computing the derivatives efficiently
When we move to the case of second-order derivatives ($g=2$), we need to distinguish between two cases: that of a partial derivative that is second order in a single variable and that of a partial derivative in two distinct variables. In the first case, we only need to elevate the derivative to the second order as follows:
@@
For the computation of the $g^{th}$ derivative, we need to perform 2h+1h+1 multiplications and one division, following Eq. 28 (or $n_{v}+1$ multiplications if $h\geq\frac{n_{v}}{2}$ according to Eq. 29 ).
@@
partial derivatives, which correspond to the sum of all possible partial derivatives at each order, with $n_{d}=\max(G_{F})$ the highest derivative order involved in the DEs, and $n_{v}$ the maximum number of variables in functions involved in the DEs (see next section for the definition of $G_{F}$). Consequently, the computational cost of computing all the partial derivatives, expressed as a function of the cost $C_{b}$ of a multiplication/division, is equal to:
@@
To compute all these partial derivatives, we only need to evaluate and multiply the value of the $d^{th}$ partial derivatives in one variable $\frac{\partial^{d}\text{Cheb}}{\partial{x_{j}}^{d}}(L_{j},x_{j})$. For each value of $d$, we have $n_{v}$ such partial derivatives to compute. In the worst case, we have to prepare all the possible partial derivatives of that form. This means that, for half of the basis states (since the other half only introduces a minus sign in front of the Chebyshev polynomials) and all sample points, in the worst case, we compute and store $2^{n-1}\times n_{v}\times n_{d}\times n_{s}$ values.